# Survey on algorithms for multi-index models

### v1

@@ Acknowledgements:
The authors would like to thank Florent Krzakala, Alex Damian, Luis Rademacher, Denny Wu and the anonymous referees for useful feedback during the completion of this work.
@@ 3 The Gaussian setting
We will first describe a general framework to estimate the support that makes strong assumptions on the input data distribution $\pi$, but in exchange enables very general choices for the link function. In particular, the Gaussian setting where $\mathbf{x}\sim\gamma_{d}$ will take centerpiece, even if some of the methods described below extend to more general settings. As discussed in Section 2.1, this guarantees the identifiability of $W:=\operatorname{ran}(L^{\scriptscriptstyle{\mathsf{T}}})$. For simplicity, we also assume $g$ is smooth (though it will be clear that only weak derivatives are needed). These assumptions enable fairly simple moment-based estimators of $W$ that are computationally tractable.
@@ 3.1 Linear estimator
indicating that $n\gg d$ samples are sufficient to obtain an accurate estimate in this setting of $\mathsf{E}[g^{\prime}(\mathbf{z})]\neq 0$. For a more precise statement we refer the reader to [38, Lemma F.11].
@@ 3.4 Application: Learning Convex Concepts
since the variance of a truncated standard normal distribution is strictly less than one. ([84] uses an inequality of [16] to establish the above inequality for general convex $K$ and to obtain quantitative bounds.) This shows that the range of $\mathsf{E}[\mathbf{y}\lparen\mathbf{x}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}-I_{d}\rparen]$ is precisely $W$. Projecting the data to the PHD subspace reduces the computational difficulty of learning convex concepts (say, using the general polynomial regression technique of [53]), because one now only has to work in $r$-dimensional space as opposed to the original $d$-dimensional space.
@@ 3.5 The Information Exponent for Single-Index Models
This analysis was put forward in the seminal works of Ben Arous, Gheissari and Jagannath [5], as well as Dudeja and Hsu [27], using slightly better algorithms. Specifically, [5] considered online SGD, which replaces the uniform gradient concentration with a sharper martingale analysis along the trajectory, leading to a slight improvement in the rate, to $O(d^{l^{\star}-1})$. On the other hand, [27] show that only two steps of a “gradient iteration” suffices:
@@ 3.7 Beyond Information and Leap Exponents via Inverse Regression
The sample complexity is thus linear in the dimension, matching the optimal information-theoretic rate. Taking for instance $g=h_{s}$ above shows an inherent advantage of this method over the previous correlation-based methods, which require $\Theta(d^{s/2})$ samples. In essence, Chen and Meka exploit the inverse regression idea, first put forward in Li’s seminal work [57]: rather than focusing solely in the regression task of modeling $\mathbf{y}|\mathbf{x}$, one can gain information by looking instead at the inverse (or ‘generative’) model $\mathbf{x}|\mathbf{y}$. The key observation is that, under Gaussian assumptions222or slightly milder; see [52] on $\mathbf{x}$, the moments $\mathsf{E}[\mathbf{x}^{\otimes k}|\mathbf{y}=y]$, seen as tensor-valued functions of $y$, are only non-constant within the index space. This idea has led to several algorithmic solutions that exploit this moment structure [24, 61, 52].
@@
Finally, let us mention that lower bounds for multi-index models against several computational classes, including SQ and differentiable queries have been recently established in [50] for product measures admitting a predefined basis aligned with the unknown subspace, and in [32] for Gaussian data for the low-degree polynomial framework. Notably, these SQ/Low-degree lower bounds identify an appropriate extension of the generative exponent to the multi-index setting — the leap generative exponent. By adapting the partial trace estimator to this setting, [32] provide an algorithm matching again this sample complexity, therefore providing a ‘complete’ statistical/computational description of the multi-index estimation problem in the high-dimensional regime (under Gaussian data).
@@ 4 Gradient span
One of the critical drawbacks of the moment methods in Section 3 is the normality assumption on $\mathbf{x}$. A standard approach to addressing this limitation is to employ score functions for non-normal distributions on $\mathbf{x}$ [e.g., 87], but otherwise the techniques are mostly unchanged. Some moment methods have also been shown to correctly estimate a subspace of the index space under weaker assumptions [e.g., 56], but these conditions are still fairly restrictive, and exhaustiveness is also only established under rather ad hoc assumptions.
@@ 4.3 Local estimates of gradients
with weights $\mathbf{w}_{i}(x)=K(\tilde{}\mathbf{x}_{i},x)$ for some kernel function $K\colon\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ (e.g., a radial kernel like $K(x,x^{\prime})=\exp(-\lVert x-x^{\prime}\rVert_{2}^{2}/h^{2})$ for a suitable bandwidth $h>0$). Related estimation strategies based on Nadaraya-Watson kernel methods, RKHS methods and finite differences were proposed and analyzed, respectively, by [76], [71], and [83]. The main downside of these approaches is that they suffer from the usual curse of dimension in nonparametric regression [81]: exponentially many (in $d$) data may be needed for accurate estimation. [76] notes, however, that estimating the EGOP may be easier than estimating the gradient field, and gives conditions under which consistency at the parametric rate may be achieved [see also 88].
@@ 5.1 Information Exponent and NNs
Expanding again this inner product in an orthonormal Hermite basis and identifying the leading order term yields [33, Lemma 1]
@@
Besides learning the hidden direction, another important question is to understand how the link function is being learnt. In [17], the authors study standard shallow NNs in the mean-field regime, and demonstrate that the network naturally learns different harmonics of the target link function at different timescales. Alternatively, [9] focuses on single-index models, and considers a ‘specialized’ shallow NN $f_{\theta}(x)=a^{\top}\sigma(Wx+b)$, with $W$ constrained to be a rank-one matrix of the form $W={\bf 1}w^{\top}$, and frozen biases drawn from a normal distribution. In other words, neurons share their input weights and only differ by a random (and frozen) bias. By adapting the analysis of [7, 27], the authors establish non-parametric rates for general link functions with a sample complexity $n=O(d^{l^{\star}})$ to learn both the index direction and the link function.

### v2

@@ Acknowledgements:
The authors would like to thank Alex Damian, Adam Klivans, Florent Krzakala, Jason Lee, Luis Rademacher, Denny Wu and the anonymous referees for useful feedback during the completion of this work.
@@ 3 The Gaussian setting
We will first describe a general framework to estimate the support that makes strong assumptions on the input data distribution $\pi$, but in exchange enables very general choices for the link function. In particular, the Gaussian setting where $\mathbf{x}\sim\gamma_{d}$ will take centerpiece, even if some of the methods described below extend to more general settings. [^2] The remarkable work of [21] uses a smoothed analysis framework to analyze algorithms for learning multi-index models under worst-case distributions on $\mathbf{x}$. In their framework, the goal of estimation is to compete against smoothed target functions. Also, the works of [60, 22] consider learning in the presence of distribution-shift, where the distribution of $\mathbf{x}$ may differ from Gaussian at “test time”. As discussed in Section 2.1, this guarantees the identifiability of $W:=\operatorname{ran}(L^{\scriptscriptstyle{\mathsf{T}}})$. For simplicity, we also assume $g$ is smooth (though it will be clear that only weak derivatives are needed). These assumptions enable fairly simple moment-based estimators of $W$ that are computationally tractable.
@@ 3.1 Linear estimator
indicating that $n\gg d$ samples are sufficient to obtain an accurate estimate in this setting of $\mathsf{E}[g^{\prime}(\mathbf{z})]\neq 0$. For a more precise statement we refer the reader to [42, Lemma F.11].
@@ 3.4 Application: Learning Convex Concepts
since the variance of a truncated standard normal distribution is strictly less than one; see [60, Appendix B and Appendix C] for quantitative bounds. This shows that the range of $\mathsf{E}[\mathbf{y}\lparen\mathbf{x}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}-I_{d}\rparen]$ is precisely $W$. Projecting the data to the PHD subspace reduces the computational difficulty of learning convex concepts (say, using the general polynomial regression technique of [57]), because one now only has to work in $r$-dimensional space as opposed to the original $d$-dimensional space.
@@ 3.5 The Information Exponent for Single-Index Models
This analysis was put forward in the seminal works of Ben Arous, Gheissari and Jagannath [5], as well as Dudeja and Hsu [29], using slightly better algorithms. Specifically, [5] considered online SGD, which replaces the uniform gradient concentration with a sharper martingale analysis along the trajectory, leading to a slight improvement in the rate, to $O(d^{l^{\star}-1})$  [^3] The cases $l^{\star}=1$ and $l^{\star}=2$ yield rates of $O(d)$ and $O(d\log d)$ respectively. This rate follows by studying the growth of a polynomial ODE of the form $\dot{m}=(1-m^{2})m^{l^{\star}-1}$, which behaves differently for $l^{\star}>2$.. On the other hand, [29] show that only two steps of a “gradient iteration” suffices:
@@ 3.7 Beyond Information and Leap Exponents via Inverse Regression
The sample complexity is thus linear in the dimension, matching the optimal information-theoretic rate. Taking for instance $g=h_{s}$ above shows an inherent advantage of this method over the previous correlation-based methods, which require $\Theta(d^{s/2})$ samples. In essence, Chen and Meka exploit the inverse regression idea, first put forward in Li’s seminal work [63]: rather than focusing solely in the regression task of modeling $\mathbf{y}|\mathbf{x}$, one can gain information by looking instead at the inverse (or ‘generative’) model $\mathbf{x}|\mathbf{y}$. The key observation is that, under Gaussian assumptions on $\mathbf{x}$, the moments $\mathsf{E}[\mathbf{x}^{\otimes k}|\mathbf{y}=y]$, seen as tensor-valued functions of $y$, are only non-constant within the index space. [^4] Similar results hold under slightly milder conditions on $\mathbf{x}$; see [56]. This idea has led to several algorithmic solutions that exploit this moment structure [26, 67, 56].
@@
Similarly, [61] recently extended the analysis of [75] to the multi-index setting, establishing weak recovery of at least one direction inside the above index space using a spectral method.
Finally, let us mention that lower bounds for multi-index models against several computational classes, including SQ and differentiable queries have been recently established in [54] for product measures admitting a predefined basis aligned with the unknown subspace, and in [36] for Gaussian data for the low-degree polynomial framework. Notably, these SQ/Low-degree lower bounds identify an appropriate extension of the generative exponent to the multi-index setting — the leap generative exponent. By adapting the partial trace estimator to this setting, [36] provide an algorithm matching again this sample complexity, therefore providing a ‘complete’ statistical/computational description of the multi-index estimation problem in the high-dimensional regime (under Gaussian data); see also [32, 33] for concurrent works that establish both an SQ lower bound, and also provide an upper bound using a subspace conditioning algorithm.
@@ 4 Gradient span
One of the critical drawbacks of the moment methods in Section 3 is the normality assumption on $\mathbf{x}$. A standard approach to addressing this limitation is to employ score functions for non-normal distributions on $\mathbf{x}$ [e.g., 93], but otherwise the techniques are mostly unchanged. Some moment methods have also been shown to correctly estimate a subspace of the index space under weaker assumptions [e.g., 62], but these conditions are still fairly restrictive, and exhaustiveness is also only established under rather ad hoc assumptions.
@@ 4.3 Local estimates of gradients
with weights $\mathbf{w}_{i}(x)=K(\tilde{}\mathbf{x}_{i},x)$ for some kernel function $K\colon\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ (e.g., a radial kernel like $K(x,x^{\prime})=\exp(-\lVert x-x^{\prime}\rVert_{2}^{2}/h^{2})$ for a suitable bandwidth $h>0$). Related estimation strategies based on Nadaraya-Watson kernel methods, RKHS methods and finite differences were proposed and analyzed, respectively, by [82], [77], and [89]. The main downside of these approaches is that they suffer from the usual curse of dimension in nonparametric regression [87]: exponentially many (in $d$) data may be needed for accurate estimation. [82] notes, however, that estimating the EGOP may be easier than estimating the gradient field, and gives conditions under which consistency at the parametric rate may be achieved [see also 94].
@@ 5.1 Information Exponent and NNs
Expanding again this inner product in an orthonormal Hermite basis and identifying the leading order term yields [37, Lemma 1]
@@
Besides learning the hidden direction, another important question is to understand how the link function is being learnt. In [16], the authors study standard shallow NNs in the mean-field regime, and demonstrate that the network naturally learns different harmonics of the target link function at different timescales. Alternatively, [9] focuses on single-index models, and considers a ‘specialized’ shallow NN $f_{\theta}(x)=a^{\top}\sigma(Wx+b)$, with $W$ constrained to be a rank-one matrix of the form $W={\bf 1}w^{\top}$, and frozen biases drawn from a normal distribution. In other words, neurons share their input weights and only differ by a random (and frozen) bias. By adapting the analysis of [7, 29], the authors establish non-parametric rates for general link functions with a sample complexity $n=O(d^{l^{\star}})$ to learn both the index direction and the link function. Remarkably, the efficient algorithm of [23] uses an ingenious enumeration technique based on lattice polynomials to learn any link function specified by a (homogeneous) deep ReLU network.